OpenStudy (anonymous):

what are 2 binomials whose product is x^2-25?? please help

OpenStudy (anonymous):

Difference of squares. \[a^2-b^2=(a+b)(a-b) \]

OpenStudy (anonymous):

Ok so what would that make the binomials?

OpenStudy (anonymous):

I just hate binomials i can't ever get them.

OpenStudy (lgbasallote):

a+ b <---first binomial a - b<---second binomial just substitue a and b to the right values

OpenStudy (anonymous):

Yes that's what I am having trouble finding though. Because the question states write 2 binomials whose product is x squared - 25

OpenStudy (anonymous):

I have no clue, i know that sounds stupid

OpenStudy (anonymous):

I just hate doing math anyways so it's confusing to me even when i do binomials.

OpenStudy (lgbasallote):

okay here's a hint \[a^2 - b^2 \implies (a+b)(a-b)\] so \[x^2 - 4 \implies (x + 2)(x-2)\] \[x^2 - 9 \implies (x+3)(x-3)\] \[x^2 - 16 \implies (x+4)(x-4)\] are you getting it now?

OpenStudy (anonymous):

so a bionomial for x squared -25 could be (x+5)?

OpenStudy (lgbasallote):

yes that's one of the binomials...what's the other?

OpenStudy (anonymous):

would it be (x-5)?

OpenStudy (lgbasallote):

yes. the binomials for x^2 - 25 is (x+5)(x-5) congrats

OpenStudy (anonymous):

well see when i typed that in my program on my computer i have to do this math on it said it was incorrect

OpenStudy (lgbasallote):

that's why i hate those computers lol. so picky

OpenStudy (lgbasallote):

try putting them in one by one

OpenStudy (anonymous):

i did i tried the x + 5 first

OpenStudy (lgbasallote):

shouldnt make a difference but...try putting x-5 first?

OpenStudy (anonymous):

Should I put the ( ) or not?

OpenStudy (anonymous):

and it says type your answer in factor form so what would that be

OpenStudy (lgbasallote):

(x+5)(x-5)

OpenStudy (lgbasallote):

^factor form

OpenStudy (anonymous):

so the (x+5)(x-5) is factor form?

OpenStudy (lgbasallote):

yes (x+5)(x-5) is factor form...

OpenStudy (anonymous):

can i ask you another math question

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

The number of ways a teacher can award different prizes to 2 students in class having n students is given by the formula p(n)=n(n-1)

OpenStudy (anonymous):

use this formula to determeine the number of ways a teacher can award different prizes to 2 students in a class having 6 students

OpenStudy (lgbasallote):

substitute 6 into n in the formula

OpenStudy (anonymous):

then it says to rewrite the formula by multiplying the factors

OpenStudy (lgbasallote):

yes. substitute 6 into n in the formula...

OpenStudy (anonymous):

ok so the formula would be P(6)=6(6-1)?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

just a couple more questions i promise

OpenStudy (anonymous):

when i tried typing that the other way it said multiply the two factors of the expression for p(n) given in the problem statement to write the expression in a different way

OpenStudy (lgbasallote):

i dont get what you're asking

OpenStudy (anonymous):

It's just saying to rewrite the formula in a different way

OpenStudy (anonymous):

by multiplying the factors

OpenStudy (lgbasallote):

just substitute 6 into n...isnt that what you did?

OpenStudy (anonymous):

Yea i did that for the first answer but it's not right for the second one

OpenStudy (lgbasallote):

now multiply

OpenStudy (lgbasallote):

simplify it

OpenStudy (anonymous):

so the answer would be 30 right that simplified

OpenStudy (anonymous):

but that's not the formula

OpenStudy (anonymous):

can i give u my math xl login info and let you look at it?

OpenStudy (lgbasallote):

the formula is what you wrote first

OpenStudy (lgbasallote):

the simplified is the second answer

OpenStudy (anonymous):

So the first formula was p(n)=n(n-1) then it was P(6)=6(6-1) then it was P(6)=6(5) and then P(6)=30

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

But when it said to multiply factors and write the formula over I have no idea

OpenStudy (lgbasallote):

multiply the factors means you multiply the factors <--that will give you 30 write the formula over is just copying the formula but substitute 6

OpenStudy (anonymous):

ok so my answer box already has P(N)= so i need to put 6 X 5

OpenStudy (lgbasallote):

if the start is p(n) = then you rewrite the formula... n(n-1)

OpenStudy (anonymous):

omg it said it wasn't in correct form that time but it was correct

OpenStudy (lgbasallote):

lol. these computers are confusing me now o.O

OpenStudy (anonymous):

I know right I don't know how to put it

OpenStudy (lgbasallote):

me neither.

OpenStudy (lgbasallote):

sorry kid. cant help you with your computer problems.

OpenStudy (lgbasallote):

i have to go now so i wish you luck that you finish this

OpenStudy (anonymous):

thanks soooo much for your help

OpenStudy (lgbasallote):

welcome